A $200,000 loan amortized over 8 years at an interest rate of 10% per year requires payments of $21,215.85 to completely remove the loan when interest is charged on the unrecovered balance of the principal. If interest is charged on the original principal instead of the unrecovered balance, what is the loan balance after 8 years provided the same $21,215.85 payments are made each year? The loan balance is $ .
Answer #1

Principal-only annual loan repayment ($) = Loan amount / P/A(10%, 8) = 200,000 / 5.3349** = 37,488.99

Annual difference in loan balance ($) = 37,488.99 - 21,215.85 = 16,273.14

Total difference in loan balance after 8 years ($) = 16,273.14 x 8 = 130,185.12